OpenStudy (anonymous):

Zero of 2 and zero of 4 having multiplicity 2; f(1) = -18

OpenStudy (anonymous):

Find a polynomial function

OpenStudy (aum):

If 2 is a zero, then (x - 2) is a factor. If 4 is a zero with multiplicity 2, then (x - 4)^2 is a factor. The polynomial is: f(x) = a * (x-2) * (x-4)^2 where a is a constant. f(1) = -18. Set x = 1. f(1) = a * (1-2) * (1-4)^2 = -18 a * (-1) * (9) = -18 -9a = -18 a = 2 f(x) = 2(x-2)(x-4)^2